(a) what are the period and speed of a person on a carousel if the person has an acceleration with a magnitude of o.96 m/s2 when she is standing 4.6 m from the axis? period speed m/s (b) What are her acceleration magnitude and speed if she then moves to a distance of 1.9 m from the carousel center and the carousel keeps rotating with the same period? m/s acceleration speed m/s
Answer #1

Given that

accelaration a=0.96 m/s^2

DISTANCE s=4.6 M

NOW WE FIND THE TIME

TIME T=[2s/a]^1/2=[2*4.6/0.96]^1/2=3.1 s

now we speed

=>speed V=[2*0.96*4.6]^1/2=2.972 m/s

.......................................................................................................................................

Given that

distance S=1.9 m

time t=3.1 s

now we find the acceleration

acceleration a=2s/t^2=2*1.9/3.1^2=0.4 m/s^2

now we find the speed

speed V=distance /time =1.9/3.1=0.613 m/s